Solve
2g + 6f + c = -10
4g - 2f + c = -5
-2g + 2g + c = -2

Answer by Alan3354(69443)

Solve
2g + 6f + c = -10
4g - 2f + c = -5
-2g + 2g + c = -2
--------------------
-2g + 2g + c = -2
c = -2
===========
2g + 6f + c = -10
4g - 2f + c = -5
c = -2
----
2g + 6f = -8 --> 4g + 12f = -16
4g - 2f = -3
4g + 12f = -16
--------------- Subtract
-14f = 13
f = -13/14
-------------
2g + 6f = -8
g - 39/14 = -4
g = -17/14